1- Mr. Nailor invests $29,000 in a money market account at his local bank. He receives annual interest of 8% for 7 years. How much return will his investment earn during this time period? Use Appendix A to calculate the answer.

  • $49,706

  • $20,706

  • $47,390

  • $16,907

2-

Carol Thomas will pay out $8,000 at the end of the year 2, $10,000 at the end of year 3, and receive $12,000 at the end of year 4. With an interest rate of 11 percent, what is the net value of the payments vs. receipts in today's dollars? Use Appendix B to calculate the answer.

  • ($5,898)

  • ($13,806)

  • ($9,112)

  • ($21,714)
Answer #1

Answer a

Future value = Present value * (1+ rate per period)^ no. of periods

= 29000*1.08^7

= 29000*1.714

= $49706

Return earned on the investment = Future value - Present value

= 49706-29000

= $20706

Note: As per Chegg guidelines, when multiple questions are posted, we are mandated do only the first question. So please post the questions individually.